OpenStudy (anonymous):

A freight train travels at 30 mph. If the freight train is 300 miles ahead of an express train traveling at 55 mph, how long will the express train take to overtake the freight train?

OpenStudy (ghazi):

A freight train travels at 30 mph. If the freight train is 300 miles ahead of an express train traveling at 55 mph, how long will the express train take to overtake the freight train? ** let x=miles freight train traveled x+300=miles express train traveled Travel time=distance/speed x/30=(x+300)/55 55x=30x+9000 25x=9000 x=360 miles Travel time=x/30=360/30=12 hrs ans: The express train will overtake the freight train in 12 hrs